# A Group Theoretic Metric for Robot State Estimation Leveraging Chebyshev Interpolation

### 30 Jan 2024 v1

@@ Abstract
We propose a new metric for robot state estimation based on the recently introduced $\text{SE}_{2}(3)$ Lie group definition. Our metric is related to prior metrics for SLAM but explicitly takes into account the linear velocity of the state estimate, improving over current pose-based trajectory analysis. This has the benefit of providing a single, quantitative metric to evaluate state estimation algorithms against, while being compatible with existing tools and libraries. Since ground truth data generally consists of pose data from motion capture systems, we also propose an approach to compute the ground truth linear velocity based on polynomial interpolation. Using Chebyshev interpolation and a pseudospectral parameterization, we can accurately estimate the ground truth linear velocity of the trajectory in an optimal fashion with best approximation error. We demonstrate how this approach performs on multiple robotic platforms where accurate state estimation is vital, and compare it to alternative approaches such as finite differences. The pseudospectral parameterization also provides a means of trajectory data compression as an additional benefit. Experimental results show our method provides a valid and accurate means of comparing state estimation systems, which is also easy to interpret and report.
@@ I INTRODUCTION
In this work, we propose a singular metric which encapsulates all elements of the generalized state vector. Our metric uses the recently proposed Lie Group $\text{SE}_{2}(3)$[5], and thus is easy to compute and add to existing software libraries. We refer to this metric as the Absolute State Error (ASE).
@@ III-A Lie Group and Lie Algebra
The special orthogonal group in 3D is the Lie group $\text{SO}(3)=\{\mathbf{R}\in\text{GL}_{3}(\mathbb{R})~{}|~{}\mathbf{R}\mathbf{R}^{T}=\mathbf{I},\text{det}\mathbf{R}=+1\}$. It represents all 3D rotation matrices, with the group operation as matrix multiplication, and the inverse is the matrix transpose. The tangent space at the identity for this manifold is the Lie algebra $\mathfrak{so}(3)$ which is given by the space of all $3\times 3$ skew symmetric matrices. The vee and hat operators are used to convert the matrix to a corresponding vector and vice-versa [33]. The exponential map associates an element of the Lie algebra to the Lie group $\mathbf{R}$: $\mathfrak{so}(3)\rightarrow\text{SO}(3)$. Conversely, the logarithm map converts the Lie group element to the corresponding Lie algebra $\text{SO}(3)\rightarrow\mathfrak{so}(3)$. For more details, we refer the reader to [34].
Similarly, the 3D special Euclidean group is given by $\text{SE}(3)=\{\mathbf{T}\in\text{GL}_{4}(\mathbb{R})~{}|~{}\mathbf{R}\in\text{\text{SO}}(3),\mathbf{p}\in\mathbb{R}^{3}\}$ and is the set of rigid transformations on $\mathbb{R}^{3}$. The group operation, inverse, exponential map and logarithm map are similar to $\text{SO}(3)$, with the addition of the translation vector as an additional element of the matrix operations [34, 33].
@@ III-B Trajectory Evaluation
Given the estimated trajectory $\mathbf{P}_{1},...,\mathbf{P}_{n}\in\text{SE}(3)$ and the true trajectory $\mathbf{Q}_{1},...,\mathbf{Q}_{n}\in\text{SE}(3)$, the Root Mean Square Error (RMSE) for RPE is given by
@@ IV-A State Definition
where $\mathbf{R}_{i}\in\text{SO}(3)$ is the rotation, $\mathbf{p}_{i}\in\mathbb{R}^{3}$ is the translation vector, and $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the linear velocity at time index $i$.
[…]
Our metric leverages the $\text{SE}_{2}(3)$ Lie group as proposed by [5]. This group is also termed as “direct spatial isometries”. Similar to how $\text{SE}(3)$ extends $\text{SO}(3)$ with the addition of the translation vector, $\text{SE}_{2}(3)$ analogously extends $\text{SE}(3)$ with the linear velocity vector.
We tweak the $\text{SE}_{2}(3)$ group definition in [5] to be equivalently defined as
@@
While we don’t explicitly use the Lie algebra aspects of the $\text{SE}_{2}(3)$ group in our metric, these properties are beneficial for other aspects, such as alignment as discussed below.
@@ IV-C Alignment
Let the similarity transformation be parameterized by $S=\{s,\mathbf{R},\mathbf{t}\}$, where $s\in\mathbb{R}$, $\mathbf{R}\in\text{SO}(3)$, and $\mathbf{t}\in\mathbb{R}^{3}$. $S$ performs the alignment transformation as
@@ IV-D Absolute State Error
We denote our proposed metric as Absolute State Error to represent the global nature of the metric as well as fact that it explicitly models state estimates. Given an estimated state trajectory $\mathbf{P}_{1:N}$ and the ground truth states as $\mathbf{Q}_{1:N}$, alongwith the alignment transformation $\mathbf{S}$ we can compute the Absolute State Error at time step i as
@@
$\text{RMSE}(\mathbf{E}_{1:N})=\Bigg{(}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{E}_{i}-I_{5\times 5}\|^{2}_{F}\Bigg{)}^{1/2}$ (8)
@@ V-A Chebyshev Series & Polynomials
A Chebyshev series is an infinite series akin to the Fourier series which forms an orthogonal basis defined on a unit circle. It is defined for functions $f(x)$ on the interval $\left[-1,1\right]$ and the series decomposition is given by
$f(t)=\sum_{k=0}^{\infty}a_{k}T_{k}(t)\;\;\mbox{with}\;\;a_{k}=\frac{2}{\pi}\int_{-1}^{1}\frac{f(t)T_{k}(t)}{\sqrt{1-t^{2}}}$ (9)
@@
Given an arbitrary real function $f$ on $\left[-1,1\right]$, we can find the best approximation to $f$ with a Chebyshev polynomial of degree $N$ (viz. Weierstrass Approximation Theorem [37]) by interpolating at the Chebyshev-Gauss-Lobatto (or simply Chebyshev) points, defined as:
@@
Interpolation at Chebyshev points ensures efficient convergence of the approximation and avoids issues such as Runge’s phenomenon [38]. Furthermore, given the interpolant values, we can make use of the Fast Fourier Transform (FFT) to efficiently obtain the coefficients $a_{k}$ of the truncated Chebyshev series. A visualization is provided in figure 2.
@@ V-B Barycentric Interpolation
Since this is a linear operation, we can reparameterize (12) as an efficient inner product $f(x)=\textbf{f}\cdot\textbf{w}$ where f is a vector of all the values of $f$ at the Chebyshev points, and w is an $(N+1)$ vector of Barycentric weights.
@@ V-C Differentiation via Interpolation
Spectral collocation methods yield an efficient process for obtaining derivatives of the approximating polynomial via the differentiation matrix. Thus, the derivatives of the interpolating polynomial can be efficiently computed via a matrix-vector product. This is useful when performing optimization as we can compute the derivatives of arbitrary functions for (almost) free.
@@ V-D Fitting to Noisy Data
The key idea is to express the least-squares polynomial fit in terms of the pseudo-spectral parameterization. Given the $m$ vector data samples $\mathbf{Z}$ at discrete time indices, we can minimize the following objective to get optimal the polynomial fit $p^{*}(\mathbf{X})$ at the pseudo-spectral points:
@@
The main mechanism used is the barycentric interpolation formula 12 which predicts the state $\mathbf{x}(i)$ at any arbitrary time $i$ as $\mathbf{x}_{i}=\mathbf{X}\cdot\mathbf{w}_{i}$, with $\Omega$ refering to the covariance of the measuring device. As a least-squares objective 13 can be minimized using non-linear programming techniques.
@@ VI COMPUTING TRUE LINEAR VELOCITY
We propose to use differentiation via interpolation to estimate the ground truth linear velocity. While most approaches compute the linear velocity using the method of finite differences, differentiation via interpolation is theoretically advantageous, as discussed in section V-C. Given the ground truth translation data from the motion capture device, we use this data as our noisy measurements in 13. By performing a nonlinear least squares optimization, we can obtain a pseudo-spectral parameterization $\mathbf{C}$ of an interpolating polynomial of degree $N$ which represents the translation data. We can use the covariance of the motion capture device as the values for $\omega$ to specify our belief in the device’s accuracy.
Finally, to compute the true linear velocity, we multiply the values of the polynomial with the $(N+1)\times(N+1)$ differentiation matrix $\mathbf{D}_{N}$, as defined in [41, p. 53]:
@@ VIII CONCLUSION
In this work, we have proposed a new metric for robot state estimation evaluation, which leverages the properties of the recently proposed $\text{SE}_{2}(3)$ Lie group. Due to its group properties, this metric explicitly takes into account the linear velocity component of the generalized state vector, while providing a singular comparative value, and is easy to incorporate into existing software libraries and pipelines.

### 25 Mar 2024 v2

@@ Abstract
We propose a new metric for robot state estimation based on the recently introduced $\mathrm{SE}_{2}(3)$ Lie group definition. Our metric is related to prior metrics for SLAM but explicitly takes into account the linear velocity of the state estimate, improving over current pose-based trajectory analysis. This has the benefit of providing a single, quantitative metric to evaluate state estimation algorithms against, while being compatible with existing tools and libraries. Since ground truth data generally consists of pose data from motion capture systems, we also propose an approach to compute the ground truth linear velocity based on polynomial interpolation. Using Chebyshev interpolation and a pseudospectral parameterization, we can accurately estimate the ground truth linear velocity of the trajectory in an optimal fashion with best approximation error. We demonstrate how this approach performs on multiple robotic platforms where accurate state estimation is vital, and compare it to alternative approaches such as finite differences. The pseudospectral parameterization also provides a means of trajectory data compression as an additional benefit. Experimental results show our method provides a valid and accurate means of comparing state estimation systems, which is also easy to interpret and report.
@@ I INTRODUCTION
In this work, we propose a singular metric which encapsulates all elements of the generalized state vector. Our metric uses the recently proposed Lie Group $\mathrm{SE}_{2}(3)$[5], and thus is easy to compute and add to existing software libraries. We refer to this metric as the Absolute State Error (ASE).
@@ III-A Lie Group and Lie Algebra
The special orthogonal group in 3D is the Lie group $\mathrm{SO}(3)=\{\mathbf{R}\in\text{GL}_{3}(\mathbb{R})~{}|~{}\mathbf{R}\mathbf{R}^{T}=\mathbf{I},\mathrm{det}\mathbf{R}=+1\}$. It represents all 3D rotation matrices, with the group operation as matrix multiplication, and the inverse is the matrix transpose. The tangent space at the identity for this manifold is the Lie algebra $\mathfrak{so}(3)$ which is given by the space of all $3\times 3$ skew symmetric matrices. The vee and hat operators are used to convert the matrix to a corresponding vector and vice-versa [33]. The exponential map associates an element of the Lie algebra to the Lie group $\mathbf{R}$: $\mathfrak{so}(3)\rightarrow\mathrm{SO}(3)$. Conversely, the logarithm map converts the Lie group element to the corresponding Lie algebra $\mathrm{SO}(3)\rightarrow\mathfrak{so}(3)$. For more details, we refer the reader to [34].
Similarly, the 3D special Euclidean group is given by $\mathrm{SE}(3)=\{\mathbf{T}\in\text{GL}_{4}(\mathbb{R})~{}|~{}\mathbf{R}\in\mathrm{SO}(3),\mathbf{p}\in\mathbb{R}^{3}\}$ and is the set of rigid transformations on $\mathbb{R}^{3}$. The group operation, inverse, exponential map and logarithm map are similar to $\mathrm{SO}(3)$, with the addition of the translation vector as an additional element of the matrix operations [34, 33].
@@ III-B Trajectory Evaluation
Given the estimated trajectory $\mathbf{P}_{1},...,\mathbf{P}_{n}\in\mathrm{SE}(3)$ and the true trajectory $\mathbf{Q}_{1},...,\mathbf{Q}_{n}\in\mathrm{SE}(3)$, the Root Mean Square Error (RMSE) for RPE is given by
@@
$\mathbf{E}_{i}=(\mathbf{Q}_{i}^{-1}\mathbf{Q}_{i+\Delta})^{-1}(\mathbf{P}_{i}^{-1}\mathbf{P}_{i+\Delta})$ (2)
@@ IV-A State Definition
where $\mathbf{R}_{i}\in\mathrm{SO}(3)$ is the rotation, $\mathbf{p}_{i}\in\mathbb{R}^{3}$ is the translation vector, and $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the linear velocity at time index $i$.
IV-B $\mathrm{SE}_{2}(3)$ Group
Our metric leverages the $\mathrm{SE}_{2}(3)$ Lie group as proposed by [5]. This group is also termed as “direct spatial isometries”. Similar to how $\mathrm{SE}(3)$ extends $\mathrm{SO}(3)$ with the addition of the translation vector, $\mathrm{SE}_{2}(3)$ analogously extends $\mathrm{SE}(3)$ with the linear velocity vector.
We tweak the $\mathrm{SE}_{2}(3)$ group definition in [5] to be equivalently defined as
@@ IV-B $\mathrm{SE}_{2}(3)$ Group
The definition of $\mathrm{SE}_{2}(3)$ thus aids us in defining closed form group operations which are also efficient to compute.
While we don’t explicitly use the Lie algebra aspects of the $\mathrm{SE}_{2}(3)$ group in our metric, these properties are beneficial for other aspects, such as alignment as discussed below.
@@ IV-C Alignment
Let the similarity transformation be parameterized by $S=\{s,\mathbf{R},\mathbf{t}\}$, where $s\in\mathbb{R}$, $\mathbf{R}\in\mathrm{SO}(3)$, and $\mathbf{t}\in\mathbb{R}^{3}$. $S$ performs the alignment transformation as
@@ IV-D Absolute State Error
We denote our proposed metric as Absolute State Error to represent the global nature of the metric as well as fact that it explicitly models state estimates. Given an estimated state trajectory $\mathbf{P}_{1:N}\in\mathrm{SE}_{2}(3)$ and the ground truth states as $\mathbf{Q}_{1:N}\in\mathrm{SE}_{2}(3)$, alongwith the alignment transformation $\mathbf{S}$ we can compute the Absolute State Error at time step i as
@@
$\text{RMSE}(\mathbf{E}_{1:N})=\Bigg{(}\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{E}_{i}-\mathbf{I}_{5\times 5}\|^{2}_{F}\Bigg{)}^{1/2}$ (10)
@@ V-A Chebyshev Series & Polynomials
A Chebyshev series is an infinite series akin to the Fourier series which forms an orthogonal basis defined on a unit circle. It is defined for functions $f(x)$ on the interval $\left[-1,1\right]$, though it can be scaled to arbitrary bounds (e.g. via min-max normalization), and the series decomposition is given by
[…]
@@
Given an arbitrary real function $f$ on $\left[-1,1\right]$, we can find the best approximation to $f$ with a Chebyshev polynomial of degree $N$. This is due to the Weierstrass Approximation Theorem [37]) which we recapitulate below
Theorem 1 (Weierstrass Approximation Theorem)
Let $f$ be a continuous function on $\left[-1,1\right]$, and let $\epsilon>0$ be arbitrary. Then there exists a polynomial $p$ such that
$\|f-p\|<\epsilon$ (13)
A generalized proof can be found in [38]. Thus, we can find the best approximation to any function by interpolating at the Chebyshev-Gauss-Lobatto (or simply Chebyshev) points, defined as:
@@
Interpolation at Chebyshev points ensures efficient convergence of the approximation and avoids issues such as Runge’s phenomenon [39] which causes undesirable oscillation of the interpolant at the endpoints. Furthermore, given the interpolant values, we can make use of the Fast Fourier Transform (FFT) to efficiently obtain the coefficients $a_{k}$ of the truncated Chebyshev series (the Chebyshev series upto a specified degree $N$). A visualization is provided in figure 2.
@@ V-B Barycentric Interpolation
An additional benefit of Barycentric Interpolation is that it yields a linear form which can be computed via matrix-vector products. Since this is a linear operation, we can reparameterize (15) as an efficient inner product $f(x)=\textbf{f}\cdot\textbf{w}$ where f is a vector of all the values of $f$ at the Chebyshev points, and w is an $(N+1)$ vector of Barycentric weights.
@@ V-C Differentiation via Interpolation
Spectral collocation methods, commonly used in the study of differential equations [42], yield an efficient process for obtaining derivatives of the approximating polynomial via the differentiation matrix. Thus, the derivatives of the interpolating polynomial can be efficiently computed via a matrix-vector product. This is useful when performing optimization as we can compute the derivatives of arbitrary functions for (almost) free.
@@ V-D Fitting to Noisy Data
The key idea is to express the least-squares polynomial fit in terms of the pseudo-spectral parameterization. Pseudo-spectral parameterization is a way of parameterizing a polynomial completely by its values at specific points, instead of with polynomial coefficients. Given the $m$ vector data samples $\mathbf{Z}$ at discrete time indices, we can minimize the following objective to get optimal the polynomial fit $p^{*}(\mathbf{X})$ at the pseudo-spectral points:
@@
The main mechanism used is the barycentric interpolation formula 15 which predicts the state $\mathbf{x}(i)$ at any arbitrary time $i$ as $\mathbf{x}_{i}=\mathbf{X}\cdot\mathbf{w}_{i}$, with $\Omega$ refering to the covariance of the measuring device. As a least-squares objective 16 can be minimized using non-linear programming techniques.
@@ VI COMPUTING TRUE LINEAR VELOCITY
We propose to use differentiation via interpolation to estimate the ground truth linear velocity. While most approaches compute the linear velocity using the method of finite differences, differentiation via interpolation is theoretically advantageous, as discussed in section V-C. Given the ground truth translation data from the motion capture device, we use this data as our noisy measurements in 16. By performing a nonlinear least squares optimization, we can obtain a pseudo-spectral parameterization $\mathbf{C}$ of an interpolating polynomial of degree $N$ which represents the translation data. We can use the covariance of the motion capture device as the values for $\omega$ to specify our belief in the device’s accuracy.
Finally, to compute the true linear velocity, we multiply the values of the polynomial with the $(N+1)\times(N+1)$ differentiation matrix $\mathbf{D}_{N}$, as defined in [43, p. 53]:
@@ VIII CONCLUSION
In this work, we have proposed a new metric for robot state estimation evaluation, which leverages the properties of the recently proposed $\mathrm{SE}_{2}(3)$ Lie group. Due to its group properties, this metric explicitly takes into account the linear velocity component of the generalized state vector, while providing a singular comparative value, and is easy to incorporate into existing software libraries and pipelines.